OpenStudy (anonymous):

90. A gas is heated from 263.0 K to 298.0 K and the volume is increased from 24.0 liters to 35.0 liters by moving a large piston within a cylinder. If the original pressure was 1.00 atm, what would the final pressure be?

OpenStudy (kendricklamar2014):

If the moles of the K is the same here is the equation. Laws of Gas Law: (v1*p1)/(n1*T1)=(v2*p2)/ (n2*T2) V equals pressure p equals pressure n equals moles T= temperature in Kelvin V1=24L V2=35 P1=1 atm P2=? T1=263.0 K T2=298 K n1=it would not matter because the amount of K would be the same. N2= It does not matter when you divide the same moles of K between n1 and n2 will be 1, so it is irrelevant for now. Renew equation to find Final P2 P2= (n2*T2*V1*P1)/ (V2*n1*T1) (1*298*24*1)/(35*1*263)=.7769 atm If you use qauntivive reasoning pressure decreases as you increase volume. Final Pressure= .7769atm

OpenStudy (anonymous):

thank u so much

OpenStudy (kendricklamar2014):

No problem :)